### v1

# The Effect of Personalization in FedProx: A Fine-grained Analysis on Statistical Accuracy and Communication Efficiency

Abstract
FedProx is a simple yet effective federated learning method that enables model personalization via regularization. Despite remarkable success in practice, a rigorous analysis of how such a regularization provably improves the statistical accuracy of each client’s local model hasn’t been fully established. Setting the regularization strength heuristically presents a risk, as an inappropriate choice may even degrade accuracy. This work fills in the gap by analyzing the effect of regularization on statistical accuracy, thereby providing a theoretical guideline for setting the regularization strength for achieving personalization. We prove that by adaptively choosing the regularization strength under different statistical heterogeneity, FedProx can consistently outperform pure local training and achieve a nearly minimax-optimal statistical rate. In addition, to shed light on resource allocation, we design an algorithm, provably showing that stronger personalization reduces communication complexity without increasing the computation cost overhead. Finally, our theory is validated on both synthetic and real-world datasets and its generalizability is verified in a non-convex setting.
Keywords: Personalized Federated Learning, Statistical Complexity, Communication Complexity, Minimax Optimality
1 Introduction
Federated Learning (FL) has emerged as an attractive framework for aggregating distributed data, enabling clients to collaboratively train a shared global model while preserving data privacy. In the currently prevalent paradigm (McMahan et al., 2017), FL is formulated as a finite sum minimization problem focusing on a single shared model. Nevertheless, it has been well recognized that one of the key challenges in FL is the statistical heterogeneity of the client datasets. As the participants collect their own local data, it often reflects client-specific characteristics and is not identically distributed. With high statistical heterogeneity, training a single model for all clients by minimizing their average in-sample loss becomes questionable.
To address this challenge, one solution is to relax the common model constraint and solve alternatively the following objective in FedProx (Li et al., 2020a):
$\min_{{{\bm{w}}^{(g)}},\{{{\bm{w}}^{(i)}}\}_{i\in[m]}}\sum_{i\in[m]}p_{i}\left({L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2}\right),$ (1)
where $L_{i}({\bm{w}}^{(i)},S_{i})$ is the empirical risk of the $i$-th client, and $S_{i}$ is the local dataset of the $i$-th client. The set $\{p_{i}\}_{i\in[m]}$ is a collection of weights for the clients satisfying $p_{i}\geq 0$ and $\sum_{i}p_{i}=1$. In (1), each client $i$ learns a local personalized model ${\bm{w}}^{(i)}$ by fitting its local data $S_{i}$, while enabling collaborative learning by shrinking all local models towards a common global one ${\bm{w}}^{(g)}$ with intensity controlled by $\lambda$. The smaller $\lambda$ is, the weaker the coupling of the local models the formulation would enforce thus the higher personalization is. Adjusting the strength of regularization controls the extent to which the ${\bm{w}}^{(i)}$’s is adapted to client $i$’s local data, thus the degree of personalization.
The promising empirical performance of FedProx in learning under heterogeneous data (Li et al., 2020a; Donevski et al., 2021; He et al., 2021; An et al., 2023) has led to a sequence of works studying the problem from a pure computation perspective, aiming at developing efficient algorithms under various problem assumptions to reach a (local) solution (Hanzely & Richtárik, 2020; T Dinh et al., 2020; Li et al., 2021; Acar et al., 2021; Yuan & Li, 2022). However, the fundamental question of how to characterize the accuracy of the models produced by (1) is largely unaddressed. As such, for a given set of clients, it is not even clear if introducing personalization will be beneficial, outperforming the baselines given by training a single global model or pure local training. In addition, communication cost in FL is a critical bottleneck in resource-constrained environments. Therefore, understanding when personalization will improve both training efficiency and statistical accuracy is of paramount importance. We address these questions by providing a fine-grained analysis of the effect of personalization in FedProx. Our contribution is summarized below:
$\bullet$ Nearly Optimal Statistical Accuracy. We analyze the statistical accuracy of the objective in (1), taking statistical heterogeneity into account. For each of the models, we establish a state-of-the-art convergence rate $\mathcal{O}(1/(mn))+\mathcal{O}(1/n\wedge R/n^{1/2})$, where $R$ characterizes the statistical heterogeneity and $n$ is the sample size on each client. The first term $\mathcal{O}(1/(mn))$ represents the sample complexity of training a single global model under homogeneity, which is attained by (1) when $R$ is close to zero, highlighting the benefit of FL. For the second term, the first component $\mathcal{O}(1/n)$ is the sample complexity of purely local training and the second component $\mathcal{O}(R/n^{1/2})$ demonstrates the flexibility of FedProx under different statistical heterogeneity, implying the statistical accuracy of FedProx never gets worse than purely local training’s. The fine-grained analysis demonstrates that with an adaptive choice of personalization degree characterized by $\lambda$, FedProx achieves a nearly minimax-optimal statistical accuracy across different levels of statistical heterogeneity.
$\bullet$ Provably Enhanced Communication Efficiency. As the objective (1) could be reformulated as a bi-level problem, we propose an algorithm for (1) with an explicit computation cost assessment. We establish the communication complexity as $\mathcal{O}(\kappa\frac{\lambda+\mu}{\lambda+L}\log\frac{1}{\varepsilon})$ and the local computation complexity in each communication round as $\mathcal{\widetilde{O}}({\frac{\lambda+L}{\lambda+\mu}})$. With higher personalization degree, communication complexity would decrease and the computation complexity per round would increase, demonstrating that there is a trade-off between computation and communication costs. Our analysis concludes that model personalization can reduce communication overhead without incurring additional computation expense.
$\bullet$ Empirical Validation. We conduct numerical studies on logistic regression with synthetic data and real-world data under our problem assumptions. The results corroborate our theoretical findings. We further test the numerical performance of (1) for training CNNs, showing empirically the results, although proved under the convexity assumption, generalize to nonconvex losses.
2 Related Works
In recent years, federated learning has become an attractive solution for training models locally on distributed clients, rather than transferring data to centralized data centers (Mammen, 2021; Wen et al., 2023; Beltrán et al., 2023). As each client generates its local data, statistical heterogeneity is common with data being non-identically distributed between clients (Li et al., 2020a; Ye et al., 2023). Given the variability of data in network, model personalization is a natural strategy used to improve learning accuracy for each client. Formulations enabling model personalization have been studied independently from multiple fields. For example, meta-learning (Chen et al., 2018; Jiang et al., 2019; Khodak et al., 2019; Fallah et al., 2020) assumes all models follow a common distribution. By minimizing the average loss, these methods aim to learn a meta-model that generalizes well to unseen new tasks. Multi-task learning methods (Agarwal et al., 2020; Liang et al., 2020) and transfer learning methods (Li et al., 2022; He et al., 2024a; b) study formulation similar to (1), but the statistical rate is established for either the average of all models or only the target model. Therefore, it is not clear if all clients can benefit from the procedure and learn a more accurate model than purely local training. Representation learning (Zhou et al., 2020; Wang et al., 2024), on the other hand, focuses on a different model structure where the local models can be represented as the composition of two parts, with one common to all clients and the other specific to each client.
In the FL community, personalized FL methods can be broadly divided into two main strands. Different from the works mentioned previously, studies here focus primarily on the properties of the iterates computed by the algorithms. One line of works (Arivazhagan et al., 2019; Liang et al., 2020; Singhal et al., 2021; Collins et al., 2021) is based on the representation learning formulation, where both the algorithmic convergence and statistical accuracy of the iterates have been investigated. Another line of works (Smith et al., 2017; Li et al., 2020a; Hanzely & Richtárik, 2020; Hanzely et al., 2020; 2023), including FedProx, achieves personalization by relaxing the requirement of learning a common global model through regularization techniques. In particular, algorithms and complexity lower bounds specific to the FedProx objective were studied in works (Li et al., 2020a; Hanzely et al., 2020; T Dinh et al., 2020; Hanzely & Richtárik, 2020; Li et al., 2021), see Table 2 for a detailed comparison. These works study the problem from a pure optimization perspective and cannot show personalization improves learning accuracy.
There are only a few recent works we are aware of studying the statistical accuracy of regularization-based PFL in the form of (1). Specifically, Cheng et al. (2023) investigates the asymptotic behavior of the FedProx estimator under an over-parameterized linear regression model; neither a finite-sample rate nor algorithm is provided. Chen et al. (2023b) studies the FedProx estimator under rather strong assumptions like bounded parameter space. By defining a new notion of algorithmic stability, they establish the non-asymptotic sample complexity for FedProx and induce minimax bound. However, even with some overly strong assumptions, their analysis cannot match the statistical lower bound in most cases.
Notation. $[n]$ denotes the set of integers $\{1,2,\cdots,n\}$ for any positive integer $n$. Let $\|\cdot{}\|_{2}$ denote the $\ell_{2}$ vector norm and the Frobenius matrix norm. For two non-negative sequences $a_{n}$, $b_{n}$, we denote $a_{n}=\mathcal{O}(b_{n})$ as $a_{n}/b_{n}\xrightarrow{}{1}$. $\tilde{\mathcal{O}}(\cdot)$ hides logarithmic factors.
3 Preliminaries
Considering there are $m$ clients, each client has a collection of $i.i.d$ data $S_{i}=\{z_{ij}\}_{j\in[n_{i}]}$ over possibly different data distributions $\mathcal{D}_{i}$. For each client, the ultimate goal is to find the solution that minimizes the local risk, i.e. ${{\bm{w}}_{\star}^{(i)}}\in\arg\min_{{\bm{w}}}\>\mathbb{E}_{z\sim\mathcal{D}_{i}}\ell({\bm{w}},z)$. Here, $\ell({\bm{w}},z)$ denotes the loss function of interest and ${{\bm{w}}_{\star}^{(i)}}\in\mathbb{R}^{d}$ is called the ground truth local model for client $i$. In this paper, we assume the model dimension $d$ is finite. Under such a setup, two commonly adopted approaches are pure local training (LocalTrain) and pure global training (GlobalTrain):
[…]
where $L_{i}({\bm{w}},S_{i}):=\sum_{j=1}^{n_{i}}\ell({\bm{w}},z_{ij})/n_{i}$ denotes the empirical risk of client $i$. In LocalTrain, each client individually minimizes its loss, whereas in GlobalTrain, all clients collaboratively minimize their average loss. Compared to LocalTrain, GlobalTrain integrates information from all clients to improve the learning performance of each client. When there are sufficient similarities between the data distribution $\mathcal{D}_{i}$ of different clients, GlobalTrain would be a better choice as more samples are included. However, if $\mathcal{D}_{i}$’s are highly unrelated, LocalTrain will be a more robust choice.
In practice, the data distribution is neither totally the same nor completely unrelated across clients. To handle the wide middle ground of the two extremes, FedProx solves the following problem
$\min_{{{\bm{w}}^{(g)}},\{{{\bm{w}}^{(i)}}\}_{i\in[m]}}\sum_{i=1}^{m}p_{i}\left({L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2}\right).$
The key feature of Fedprox is the incorporation of a proximal (regularization) term, which controls the personalization degree. As $\lambda\rightarrow 0$, the FedProx objective increasingly behaves like LocalTrain with higher personalization degree and reduces to $m$ separate LocalTrain problems in the end. On the other hand, as $\lambda\rightarrow\infty$, the local models ${\bm{w}}^{(i)}$ are shrunk towards the global model ${\bm{w}}^{(g)}$ with lower personalization degree and eventually degenerate to optimize a single shared model. Thus, by adjusting $\lambda$, FedProx interpolates between LocalTrain and GlobalTrain, allowing for a flexible choice of personalization degree to achieve higher modeling accuracy. It’s noticeable that this choice of $\lambda$ also affects the optimization process of FedProx. With larger $\lambda$, more frequent communication between clients is expected to ensure faster convergence, as local updates are more tightly coupled to the global model. In contrast, smaller $\lambda$ value allows for more independent updates, reducing the need for communication. In the next few sections, we provide a theoretical guarantee for the intuition provided above.
In Section 4, we analyze when the personalization level, characterized by $\lambda$, would improve the statistical accuracy of both the global model ${\bm{w}}^{(g)}$ and the local models $\{{{\bm{w}}^{(i)}}\}_{i\in[m]}$. This is in sharp contrast to most of the existing works where the error is established for the sum of them. In Section 5, we explore how such a degree of personalization will affect communication complexity.
4 When will Personalization Improve Statistical Accuracy
In this section, we analyze how the choice of $\lambda$, or the degree of personalization, would improve the statistical accuracy of the solution obtained from the FedProx problem (1). Following the existing literature (Li et al., 2023; Chen et al., 2023a; b; Duan & Wang, 2023), we define the statistical heterogeneity as the Euclidean distance between the true local models ${{\bm{w}}_{\star}^{(i)}}$ and the true global model ${{\bm{w}}_{\star}^{(g)}}:=\sum_{i=1}^{m}p_{i}{{\bm{w}}_{\star}^{(i)}}$. Specifically, we adopt the following assumption:
Assumption 1 (Statistical Heterogeneity).
There exists a nonnegative constant $R$ such that we have $\|{{\bm{w}}_{\star}^{(i)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq R^{2},\forall i\in[m]$.
Intuitively, a larger $R$ indicates a greater divergence between the clients’ data distributions, thus stronger statistical heterogeneity, and vice versa. Therefore, in the FedProx objective, the choice of $\lambda$ should be adjusted according to $R$. Next, we introduce two standard assumptions.
Assumption 2 (Strong Convexity).
The loss function $\ell(\cdot,z)$ is $\mu$-strongly convex for each local sample average, i.e. $\sum_{j\in[n_{i}]}\ell(x,z_{ij})/n_{i}\geq\sum_{j\in[n_{i}]}\ell(y,z_{ij})/n_{i}+\langle\sum_{j\in[n_{i}]}\nabla\ell(y,z_{ij})/n_{i},y-x\rangle+\frac{\mu}{2}\|x-y\|^{2},\quad\forall x,y\in\mathbb{R}^{d},i\in[m].$
Assumption 3 (Bound Gradient Variance at Optimum).
There exists a nonnegative constant $\rho$ such that $\forall i\in[m]$, we have $\mathbb{E}_{z\sim\mathcal{D}_{i}}\|\nabla\ell({{\bm{w}}_{\star}^{(i)}},z)\|^{2}\leq\rho^{2}$.
Here, the strongly convex assumption is used to establish the estimation error and is widely adopted in the statistical analysis of regularization-based personalized federated learning (T Dinh et al., 2020; Deng et al., 2020; Hanzely & Richtárik, 2020; Hanzely et al., 2020) and statistical analysis in federated learning (Chen et al., 2023b; Cheng et al., 2023). Assumption 3 is used to quantify the observation noise, and is standard in the literature (Li et al., 2020b; Chen et al., 2023b).
Next, we establish a statistical rate of the optimal solution to the Fedprox problem (1) with a fine-grained analysis. To facilitate the analysis, we first define some notations. Define $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ as the optimal solution for local models. A check on the optimality condition implies that the optimal solution for the global model is $\widetilde{\bm{w}}^{(g)}=\sum^{m}_{i=1}p_{i}\widetilde{\bm{w}}^{(i)}$. In this paper, we use global statistical error $\mathbb{E}\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and local statistical error $\mathbb{E}\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$ as the measure of statstical accuracy. Here the expectation is taken with respect to all data. Note that some existing literature adopts alternative metrics, such as average excess risk and individual excess risk (Chen et al., 2023b) to measure the statistical error. They are equivalent to our metrics under the assumptions used in this paper. We first provide a lemma to cast light on the effect of $\lambda$ on the statistical error. See Appendix A.2 for the proof.
Lemma 1.
Suppose Assumptions 1 and 2 hold. We have
$\underbrace{\sum_{i\in[m],j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left(\mu\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}+\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\rangle\right)}_{\text{model estimation error}}\leq\frac{\lambda}{2}\underbrace{(R^{2}-\sum_{i\in[m]}p_{i}\|\widetilde{\bm{w}}^{(i)}-\widetilde{\bm{w}}^{(g)}\|^{2})}_{\text{heterogeneity estimation error}}.$ (3)
Lemma 1 establishes a relationship between the local model estimation error and the statistical heterogeneity. On the right-hand side of (3), the first term $R$, which is larger than $\|{{\bm{w}}_{\star}^{(i)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$, measures the statistical heterogeneity; the second term, $\sum_{i=1}^{m}p_{i}\|\widetilde{\bm{w}}^{(i)}-\widetilde{\bm{w}}^{(g)}\|^{2}$, can be understood as the estimator of such heterogeneity. Lemma 1 implies that to achieve a tight bound on the estimation error, we should choose a $\lambda$ that “minimizes” the right-hand side of (3). When $R$ is large, one straightforward choice is $\lambda=\mathcal{O}(1/(nR^{2}))$, which will be shown to yield an error bound of order $\mathcal{O}(1/n)$, matching the performance of purely local training and is independent of $R$. When $R$ is small, $\lambda$ should be carefully chosen to obtain the full sample complexity $\mathcal{O}(1/(mn))$. In the following theorem, we provide such a choice and establish a state-of-the-art statistical rate.
Theorem 1.
Suppose Assumptions 1, 2 and 3 hold and $n_{i}=n,\forall i\in[m]$. Consider the FedProx objective in (1) with $p_{i}=\frac{1}{m},\forall i\in[m]$. If $\lambda=\frac{\rho}{\sqrt{n}R}\mathbb{I}\{R\leq n^{-1/2}\}+\frac{\rho^{2}}{nR^{2}}\mathbb{I}\{R>n^{-1/2}\}$, the global model $\widetilde{\bm{w}}^{(g)}$ and local models $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ obtained from solving the objective (1) satisfy
$\mathbb{E}\left\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\right\|^{2}\leq\frac{C_{1}}{N}+C_{2}\left(\frac{1}{n}\wedge\frac{R}{\sqrt{n}}\right),$ (4)
$\mathbb{E}\left\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\leq\frac{C_{3}}{N}+C_{4}(\frac{1}{n}\wedge\frac{R}{\sqrt{n}})+16(\frac{\rho^{2}}{\mu^{2}n}\wedge R^{2})\quad\text{ for all }i\in[m],$ (5)
where $N=mn$ and
$\begin{split}C_{1}&=\frac{\rho^{2}}{\mu^{2}}=\frac{C_{3}}{8},C_{2}=\big{(}\frac{(1+\sqrt{1+2\mu})\rho}{2\mu}\big{)}^{2}\vee\frac{2\rho}{\mu},C_{4}=\big{(}\frac{\rho(1+(1+4\mu(C_{2}+1))^{\frac{1}{2}})}{2\mu}\big{)}^{2}\wedge(8C_{2}+\frac{2\rho}{\mu}).\end{split}$ (6)
Proof.
See Appendix A.2 for (4) and Appendix A.3 for (5). ∎
Theorem 1 shows that with an appropriate choice of $\lambda$, both the global and local models achieve a fast statistical rate across a wide range of $R$. To interpret the result, first we analyze the error bound provided in (4) and (5). The two error bounds have the same order and can be decomposed into two key components. The first term, of order $\mathcal{O}(N^{-1})$, captures the benefit of collaborative learning, as $N=mn$ is the total sample size across $m$ clients. This term reflects the sample efficiency gained through collaborative learning. The second term, of order $\mathcal{O}(n^{-1}\wedge n^{-1/2}R)$, measures the cost of statistical heterogeneity. As $R$ increases, the convergence rate gradually degrades, transitioning from $\mathcal{O}(N^{-1})$ to $\mathcal{O}(n^{-1})$, showing that the error bound of FedProx interpolates between the two extremes: LocalTrain and GlobalTrain.
Next, to better understand how such an interpolation is achieved through an adaptive personalization degree, we may examine the two extreme cases: as the statistical heterogeneity $R\to\infty$, and $\lambda\to 0$, leading to a high degree of personalization. The statistical error bound converges to $\mathcal{O}(n^{-1})$, matching the rate of LocalTrain; as the statistical heterogeneity $R\to 0$, and $\lambda\to\infty$, leading to a low degree of personalization. The statistical error bound converges to $\mathcal{O}(N^{-1})$, matching the rate of GlobalTrain. As $R\in(0,\infty)$, $\lambda$ is adjusted correspondingly, making FedProx always perform no worse than LocalTrain while benefiting from global training when clients share similarities.
Table 1: Results Comparison. $D:=\sup_{\bm{w},\bm{w}^{\prime}\in\mathcal{D}}\|\bm{w}-\bm{w}^{\prime}\|$ is diameter of the parameter space ($D\geq R$) and $\|\ell\|_{\infty}=\inf\{M\in\mathbb{R}:\ell(\cdot,z)\leq M,\text{for any }z\}$ is an uniform upper bound on the loss function. Statistical rate refers to $\mathbb{E}\|{{\bm{w}}_{\star}^{(i)}}-\widetilde{\bm{w}}^{(i)}\|^{2}$, with constants neglected.
Source Assumption Paradigm Statistical Rate
$R>\frac{1}{\sqrt{n}}$ $\frac{1}{m\sqrt{n}}<R\leq\frac{1}{\sqrt{n}}$ $R\leq\frac{1}{m\sqrt{n}}$
(Chen et al., 2023b) A1,2,3, $D\vee\|\ell\|_{\infty}\leq C$ Two-stage $\frac{D^{2}+\|\ell\|_{\infty}}{n}$ 1 $\frac{\|\ell\|_{\infty}}{\sqrt{n}}R$ $\frac{\|\ell\|_{\infty}}{\sqrt{m}n}$
Ours A1,2,3 One-stage $\frac{1}{n}$ $\frac{1}{\sqrt{n}}R$ $\frac{1}{mn}$
Lower Bound - - $\frac{1}{n}$ $R^{2}\vee\frac{1}{mn}$ $\frac{1}{mn}$
We now compare our theoretical results with the result up to date (Chen et al., 2023b), as summarized in Table 1. Our contribution can be summarized into three aspects:
$\bullet$ Our results are derived under more realistic assumptions. Chen et al. (2023b) imposes additional stringent conditions including a bounded parameter space $D$ and a uniform bound on the loss function $\ell(\cdot,z)$. Such conditions are difficult to satisfy in practice. For example, even a simple linear regression model with sub-Gaussian covariates would violate these assumptions.
$\bullet$ Our result is established for both global and local models with the same choice of $\lambda$. In contrast, Chen et al. (2023b) proposes a two-stage process where the global model is first estimated, followed by an additional local training phase with a different choice of $\lambda$. This two-step approach is likely the artifact of their analysis, as existing work on FedProx also shows that a single-stage optimization with a chosen $\lambda$ suffices to achieve the desired performance (T Dinh et al., 2020).
$\bullet$ We establish a state-of-the-art nearly minimax statistical rate. As shown in Table 1, in a heterogeneous case when $R\geq n^{-1/2}$, Chen et al. (2023b) establishes a rate that is at least of the order $\mathcal{O}\left(n^{-1}R^{2}\right)$ as $D\geq R$. This bound becomes increasingly loose with larger $R$. In contrast, our bound is $\mathcal{O}\left(n^{-1}\right)$, which shows that properly-tuned FedProx is always no worse than LocalTrain, independent of $R$. In a homogeneous case when $R\leq m^{-1}n^{-1/2}$, they can only establish a rate slower than $\mathcal{O}\left(1/(\sqrt{m}n)\right)$, while our result achieves a rate of $\mathcal{O}\left(1/(mn)\right)$, leveraging all $mn$ samples and matching the rate of GlobalTrain on the IID data. Compared to the established lower bound (Chen et al., 2018), our result is nearly minimax optimal. Specifically, when $R>n^{-1/2}$ and $R\leq m^{-1}n^{-1/2}$, our result matches the lower bound. There is a small gap exists in the case when $m^{-1}n^{-1/2}<R\leq n^{-1/2}$, where we achieve a rate of $\mathcal{O}(n^{-1/2}R)$ whereas the lower bound is $\mathcal{O}(R^{2})$. Such a gap is due to the cost of estimating the model difference $\{{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\}_{i=1}^{m}$ when $\lambda$ is finite (c.f. Lemma 1). The improvement of this gap is non-trivial, which we defer to future work.
5 When will Personalization Improve Communication Efficiency
Notice that if we define the local objective of each client $i$ as $h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}):={L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2}$, then the FedProx problem given by (1) can be rewritten in the following bilevel (iterated minimization) form:
$\min_{{\bm{w}}^{(g)}}F({\bm{w}}^{(g)}):=\frac{1}{m}\sum_{i=1}^{m}F_{i}({\bm{w}}^{(g)}),\quad\text{where}\quad F_{i}({\bm{w}}^{(g)}):=\min_{{{\bm{w}}^{(i)}}}~{}h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}).$ (7)
The reformulation (7) is a finite-sum minimization problem, with each component $F_{i}$ being the Moreau envelope (Moreau, 1965; Yosida, 1964) of $L_{i}$. Let ${\bm{w}}_{\star}^{(i)}({\bm{w}}^{(g)})$ be the minimizer of $h_{i}(\,\cdot\,,{\bm{w}}^{(g)})$, i.e.,
${\bm{w}}_{\star}^{(i)}({\bm{w}}^{(g)}):=\operatorname{prox}_{L_{i}/\lambda}({\bm{w}}^{(g)})=\operatorname*{arg\,min}_{{{\bm{w}}^{(i)}}}~{}h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}).$ (8)
As a standard argument about the convergence rate (T Dinh et al., 2020; Li et al., 2021; Chen et al., 2023b), we assume the loss function $\ell$ is $L$-smooth.
Assumption 4 (Smoothness).
The loss function $\ell(\cdot,z)$ is L-smooth, i.e. for any $x,y\in\mathbb{R}^{d}$
$\|\nabla\ell(x,z)-\nabla\ell(y,z)\|\leq L\|x-y\|,\quad\forall z.$ (9)
Based on Assumption 4 and the new notations in (7) and (8), the following lemma describes the optimizing properties for $F_{i}$, $h_{i}$ and ${\bm{w}}_{\star}^{(i)}$. The proof can be found in Appendix B.2.
Lemma 2.
Under Assumption 2 and 4, $F_{i}$ is $\mu_{g}$-strongly convex and $L_{g}$-smooth, with $\mu_{g}=\frac{\lambda\mu}{\lambda+\mu}$ and $L_{g}=\frac{\lambda L}{\lambda+L}$, each $h_{i}$ is $\mu_{\ell}$-strongly convex and $L_{\ell}$-smooth, with $\mu_{\ell}=\mu+\lambda$ and $L_{\ell}=L+\lambda$, and the mapping $\bm{w}_{\star}^{(i)}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is $L_{w}$-Lipschitz with $L_{w}=\frac{\lambda}{\lambda+\mu}$.
Next, we recall the properties of the Moreau envelope to guide the algorithm design.
Lemma 3 (Lemaréchal & Sagastizábal (1997)).
Under Assumption 2, each $F_{i}:\mathbb{R}^{d}\to\mathbb{R}$ is continuously differentiable, and the gradient is given by $\nabla F_{i}({\bm{w}}^{(g)})=\lambda({\bm{w}}^{(g)}-{\bm{w}}_{\star}^{(i)}({\bm{w}}^{(g)})).$
From Lemma 2 and Lemma 3, it suggests applying a simple gradient algorithm to optimize ${\bm{w}}^{(g)}$:
$\begin{split}{\bm{w}}_{t+1}^{(g)}&={\bm{w}}_{t}^{(g)}-\gamma\cdot\nabla F({\bm{w}}_{t}^{(g)})={\bm{w}}_{t}^{(g)}-\gamma\lambda\cdot\frac{1}{m}\sum_{i=1}^{m}\left({\bm{w}}_{t}^{(g)}-{\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})\right),\end{split}$ (10)
where $\bm{w}_{t}^{(g)}$ is the update at the communication round $t$, $\gamma>0$ is a step size to be properly set (cf. Theorem 2). To implement (10), in each communication round $t$ the server broadcasts ${\bm{w}}_{t}^{(g)}$ to all clients, then each client $i$ updates its local model ${\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})$ and upload to the server.
Note that executing (10) requires each client $i$ computing the minimizer ${\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})$. In general, the subproblem (8) does not have a closed-form solution. Therefore, computing the exact gradient $\nabla F$ will incur a high computation cost as well as high latency for the learning process. Leveraging recent advancements in bilevel optimization algorithm design (Ji et al., 2022), we address this issue by approximating ${\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})$ with a finite number of $K$ gradient steps. Specifically, we let each client $i$ maintain a local model ${\bm{w}}_{t,k}^{(i)}$. Per communication round $t$, ${\bm{w}}_{t,k}^{(i)}$ is initialized to be ${\bm{w}}_{t,0}^{(i)}={\bm{w}}_{t-1,K}^{(i)}$ as a warm start, and updated according to
$\begin{split}{\bm{w}}_{t,k+1}^{(i)}={\bm{w}}_{t,k}^{(i)}-\eta\nabla h_{i}({\bm{w}}_{t,k}^{(i)},{\bm{w}}_{t}^{(g)}),\quad\forall\,k=0,\ldots,K-1,\end{split}$ (11)
where $\nabla h_{i}({\bm{w}}_{t,k}^{(i)},{\bm{w}}_{t}^{(g)})$, for notation simplicity, denotes the partial gradient of $h_{i}$ with respect to ${\bm{w}}^{(i)}$. The overall procedure is summarized in Alg. 1.
Input: Initial global model ${\bm{w}}_{1}^{(g)}$, initial local models $\{{\bm{w}}_{0,K}^{(i)}\}_{i\in[m]}$, global rounds $T$, global step sizes $\gamma$, local rounds $K$, local step sizes $\eta$
Output: Local models $\{{\bm{w}}_{T,K}^{(i)}\}_{i\in[m]}$ and global model ${\bm{w}}_{T}^{(g)}$
for $t=1,\dots,T$ do
       The server sends ${\bm{w}}_{t}^{(g)}$ to client $i$, $\forall i\in[m]$
       Set ${\bm{w}}_{t,0}^{(i)}={\bm{w}}_{t-1,K}^{(i)}$;
       for $k=0,\dots,K-1$ do
             ${\bm{w}}_{t,k+1}^{(i)}={\bm{w}}_{t,k}^{(i)}-\frac{\eta}{n_{i}}\sum_{j\in[n_{i}]}\left\{\nabla\ell({\bm{w}}_{t,k}^{(i)},z_{i,j})+\lambda\left({\bm{w}}_{t,k}^{(i)}-\bm{w}_{t}^{(g)}\right)\right\}$
      Push ^$\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})=\lambda(\bm{w}_{t}^{(g)}-{\bm{w}}_{t,K}^{(i)})$ to the server;
      $\bm{w}_{t+1}^{(g)}\leftarrow\bm{w}_{t}^{(g)}-\frac{\gamma}{m}\sum_{i\in[m]}\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})$;
Algorithm 1 One-stage FedProx
Remark 1.
Alg. 1 is a customization of the general bi-level problem (Ji et al., 2022). While the algorithm design is not new, our contribution lies in analyzing how the personalization degree influences the trade-off between communication and local computation leveraging the FedProx problem structure. For simplicity, we focus on the full gradient update in our analysis; both the algorithm and the analysis can be extended to the stochastic setting with mini-batch gradients, with partial participation and incorporating variance reduction techniques.
Theorem 2.
Suppose Assumptions 2 and 4 hold. Let $\{{\bm{w}}_{t}^{(g)}\}_{t\geq 0}$ and $\{{\bm{w}}_{t,K}^{(i)}\}_{t\geq 0}$ be the sequence generated by Algorithm 1 with $\gamma<1/L_{g}$, $\eta\leq 1/L_{\ell}$, and the inner loop iteration number satisfying
$\Big{(}2+64L_{w}^{2}(1/\mu_{g})^{2}\lambda^{2}\Big{)}(1-\eta\mu_{\ell})^{K}\leq(1-\gamma L_{g})^{4},$ (12)
then ${\bm{w}}_{t}^{(g)}$ converges to $\widetilde{\bm{w}}^{(g)}$ linearly at rate $21-(\gamma\mu_{g})/2-(\gamma\mu_{g})^{2}/$ 2, and ${\bm{w}}_{t,0}^{(i)}$ converges to $\widetilde{\bm{w}}^{(i)}$ linearly at the same rate for any $i\in[m]$.
Proof.
See Appendix B.3 and B.5. ∎
Theorem 2 shows under the step size constraint $\gamma<1/L_{g}$, ${\bm{w}}_{t}^{(g)}$ converges to $\tilde{{\bm{w}}}^{(g)}$ at the same rate as the iteration (10) with exact gradient computation. Furthermore, the larger $\gamma$ is, the larger $K$ should be chosen so as to satisfy (12). This indicates a tradeoff between communication and computation: the fewer communication rounds are performed, then the more local computation is required for the algorithm to converge.
To better interpret the result and obtain insights on the influence of $\lambda$ on the convergence, we provide the communication and computation complexity under specific choices the algorithm’s tuning parameters, as described in the following Corollary 1. See Appendix B.4 for the proof.
Corollary 1.
In the same setting as Theorem 2, if choosing the step size $\eta=1/L_{\ell}=(\lambda+L)^{-1}$, $\gamma=1/(2L_{g})=\frac{\lambda+L}{2\lambda L}$, and the inner loop iteration number $K=\tilde{\mathcal{O}}\left(\frac{\lambda+L}{\lambda+\mu}\right)$, then the communication complexity and computation complexity for Algorithm 1 to find an $\varepsilon$-solution (i.e., $\|{\bm{w}}_{T}^{(g)}-\hat{{\bm{w}}}^{(g)}\|^{2}\leq\varepsilon$ and $\|{\bm{w}}_{T,0}^{(i)}-\hat{{\bm{w}}}^{(i)}\|^{2}\leq\varepsilon$) are as follows ($\kappa:=L/\mu$):
(i) # communication $\begin{split}&\text{(i) \# communication rounds}=\mathcal{O}\big{(}\kappa_{g}\log(1/\varepsilon)\big{)}=\mathcal{O}\left(\kappa\cdot\frac{\lambda+\mu}{\lambda+L}\cdot\log\frac{1}{\varepsilon}\right),\end{split}$ (13)
(ii) # gradient $\begin{split}&\text{(ii) \# gradient evaluations}=\tilde{\mathcal{O}}\left(\kappa\cdot\log\frac{1}{\varepsilon}\right).\end{split}$ (14)
$\bullet$ Communication cost. Corollary 1 shows as $\lambda$ increases from 00 to $\infty$, the communication complexity increases from $\tilde{\mathcal{O}}(1)$ to $\tilde{\mathcal{O}}(\kappa)$. The result reveals a higher personalization degree requires less communication cost to collaboratively learn the models. In the full personalization limit with $\lambda\rightarrow 0$, the order of the communication cost reduces to near constant $\tilde{\mathcal{O}}(1)$, which corresponds to pure local training. It is worth noting that even personalization can reduce communication complexity, as our ultimate goal is to accurately estimate each model so we can not simply pursue higher communication efficiency.
$\bullet$ Computation cost. In contrast to the communication complexity, the gradient evaluations per round $K$ decrease with $\lambda$ increasing. Therefore, a higher degree of personalization requires a more intensive update of the local models once receiving a new global model ${\bm{w}}^{(g)}$, so as to adapt to each client’s local data set. Note that although a small $\lambda$ requires a larger $K$, the overall computation cost given by (LABEL:eq:comp-cost) is independent of $\lambda$. As such, model personalization (small $\lambda$) can provably reduce the communication cost without any extra computation overhead.
[b] Methods Convexity Bounded Domain Computation Cost per Round Communication Cost FedProx (Li et al., 2020a) cvx (1) $\mathcal{O}(\Delta/(\rho\varepsilon))$ pFedMe (T Dinh et al., 2020) s-cvx (2) $\mathcal{O}(1/{(mR\varepsilon)})$ Ditto (Li et al., 2021) s-cvx $\mathcal{O}(1)$ $\mathcal{O}(1/\varepsilon)$ L2GD (Hanzely & Richtárik, 2020) s-cvx $\mathcal{O}(1+\frac{L}{\lambda})$ $\mathcal{O}({\frac{2\lambda\kappa}{\lambda+L}\log(\frac{1}{\varepsilon})})$ Algorithm 3 (Chen et al., 2023b) s-cvx $\mathcal{O}(\lambda/\varepsilon)$ $\mathcal{O}((\lambda\vee 1)/\varepsilon)$ Our Result s-cvx $\mathcal{O}(\frac{\lambda+L}{\lambda+\mu})$ $\mathcal{O}(\kappa\frac{\lambda+\mu}{\lambda+L}\log{\frac{1}{\varepsilon}})$
Table 2: Comparison of the communication and computation complexity in existing works of analyzing the objective in (1) (cvx denotes convexity and s-cvx denotes strong convexity).
  • (1)
    Controlled by the precision of inexact solution. $\rho>0$ measures the subproblem solution accuracy.
  • (2)
    Related to the subiteration number $R$ and the precision $v$.
We end this section with a comparison between our results and the existing algorithms for solving FedProx, as reported in Table 2. Our result shows how the personalization degree trades off the communication cost and local computation cost, whereas most of the existing works either don’t explicitly analyze the local computational cost (Li et al., 2020a; T Dinh et al., 2020), or fail to demonstrate such a trade-off (Chen et al., 2023b). Hanzely & Richtárik (2020) also reveals similar tradeoffs, but their work is a randomized method that requires more coordination between server and clients, and more importantly, when $\lambda$ is small, the computation cost per round is unbounded.
6 Simulation
In this section, we provide empirical validation for our theoretical results, evaluating FedProx on both synthetic and real datasets with convex and non-convex loss functions. We first examine the effect of personalization on statistical accuracy under different levels of heterogeneity (Section 6.2), followed by an analysis of its impact on communication efficiency (Section 6.3). Detailed descriptions of the experimental setup are available in Section 6.1, and the complete anonymized codebase is accessible at https://anonymous.4open.science/r/fedprox-bilevel.
6.1 Experimental Details
Synthetic Dataset. To generate the synthetic data, we follow a similar procedure to prior works (Li et al., 2020a) but with some modifications to align with the setup in this paper. Specifically, for each client we generate samples $(\bm{X}_{k},\bm{y}_{k})$, where the labels $\bm{y}_{k}$ are produced by a logistic regression model $\bm{y}_{k}=\text{argmax}(\sigma({\bm{w}}_{k}^{\top}\bm{X}_{k}))$, with $\sigma$ being the sigmoid function. The feature vectors $\bm{X}_{k}$ are drawn from a multivariate normal distribution $\mathcal{N}(\bm{v}_{k},\bm{\Sigma})$, where $v_{k}\sim\mathcal{N}(0,1)$ and the covariance matrix $\bm{\Sigma}$ follows a diagonal structure $\bm{\Sigma}_{j,j}=j^{-1.2}$. The heterogeneity is introduced by sampling the model weights ${\bm{w}}_{k}$ for each client from a normal distribution $\mathcal{N}(0,R)$, where $R\in[0,3]$ controls the degree of variation (statistical heterogeneity) across clients’ data distributions.
Real Dataset. We use the MNIST dataset for the real dataset analysis. We first formulate the problem as a convex classification task using multinomial logistic regression, then test the generalization of our theory to a more general, non-convex scenario by conducting experiments using a CNN classifier. Following Li et al. (2020a), to impose statistical heterogeneity, we distribute the MNIST data across clients in the way that each client only has access to $s\in\{2,4,6,8,10\}$ classes. The fewer classes each client has access to, the higher the statistical heterogeneity. To be consistent with our theory, for both the synthetic and real datasets, we sample 2,000 samples and distribute them evenly among 10 clients, and all clients fully participate in each communication round. Additional results with 100 and 1,000 clients under partial participation are provided in Appendix C.
Selection of $\lambda$. For the logistic regression on the synthetic dataset, we selected three values for $\lambda$: small $\lambda=0.02$, medium $\lambda=0.1$, and large $\lambda=0.5$. For logistic regression on the MNIST dataset, we used small $\lambda=0.5$, medium $\lambda=1.5$, and large $\lambda=2.5$. For CNN on the MNIST dataset, we used small $\lambda=0.2$, medium $\lambda=0.5$, and large $\lambda=2.5$. Such a choice is to make sure that each line in the figure is clearly separable. Additionally, we implemented an adaptive choice of $\lambda=\frac{\rho}{\sqrt{n}R}\mathbb{I}\{R\leq n^{-1/2}\}+\frac{\rho^{2}}{nR^{2}}\mathbb{I}\{R>n^{-1/2}\}$ following the formula outlined in Theorem 1, with $\rho$ set to 2. The value of $\rho$ was determined through grid search over the range $[1,10]$ to ensure optimal performance.
Selection of Step Size. For the synthetic dataset, aligned with Theorem 2, we set the global step size $\gamma=\frac{\lambda+L}{\lambda L}$ and the local step size $\eta=\frac{1}{L+\lambda}$. The smoothness constant $L$ is computed as the upper bound of the $L$-smoothness constant for logistic regression, specifically $4^{-1}\Lambda_{\text{max}}$, where $\Lambda_{\text{max}}$ denotes the largest eigenvalue of the Gram matrix $\bm{X}^{\top}\bm{X}$. The local step size for the synthetic dataset was further tuned through grid search in $\{0.001,0.01,0.1\}$. For the MNIST dataset, the global learning rate was set to $\frac{1}{\lambda}$, while the local learning rate was set to 0.01, chosen via grid search. For both GlobalTrain and LocalTrain training baselines, the learning rate was also chosen via grid search within the range $\{0.001,0.01,0.1\}$. LocalTrain is implemented using a local (S)GD, and GlobalTrain is implemented using FedAvg (McMahan et al., 2017).
Selection of Hyperparameters. For the MNIST dataset, we implemented logistic regression using the SGD solver with 5 epochs, a batch size of 32, and 20 total runs. For the CNN model on the MNIST dataset, the same hyperparameter setup was used, except the batch size was reduced to 16 to ensure stable convergence.
Optimizer. For the synthetic dataset, following the analysis in the paper, the loss function is optimized using gradient descent (GD). For the MNIST dataset, all models are implemented in PyTorch, and optimization is performed using the stochastic gradient descent (SGD) solver.
Evaluation Criterion. In terms of evaluation, for the synthetic dataset, we track the ground-truth global model ${{\bm{w}}_{\star}^{(g)}}$ and local model ${{\bm{w}}_{\star}^{(i)}}$. We compute the global statistical error as $\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and the local statistical error as $\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$. To capture the total error at the $t$-th iteration, we report both the global total error $\|\bm{w}_{t}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and the local total error $\|{\bm{w}}_{t,0}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$, which combine both optimization and statistical errors. For the MNIST dataset, we report test accuracy on the MNIST testing set as a measure of the statistical error, and training loss for each client’s training set as total error at the $t$-th iteration.
Figure 1: Statistical and total error of different methods. In the top row, subfigures (a), (b), and (c) represent statistical error, showing the difference between the minimizer and the underlying true model for Synthetic-Logit (logistic regression on synthetic data), MNIST-Logit (logistic regression on MNIST), and MNIST-CNN (CNN on MNIST), respectively. In the bottom row, subfigures (d), (e), and (f) represent total error, showing the error between the model at each round and the true model for the same datasets. (1) FedProx adapts between GlobalTrain and LocalTrain, achieving low error across varying levels of heterogeneity with a properly chosen $\lambda$; (2) Larger $\lambda$ (less personalization) requires more communication to converge, and vice versa.
6.2 Statistical Aspect: Balancing Local and Global Learning
We now explore how FedProx balances between GlobalTrain and LocalTrain by achieving different levels of personalization. For comparison, in addition to implementing FedProx with different choices of $\lambda$, we also implemented LocalTrain and GlobalTrain, with details specified in Appendx LABEL:sec:additional_exp. We run all methods to a stable convergence and measure the statistical error. The top row of Figure 1 shows the performance of these methods under different impact of statistical heterogeneity.
From Figure 1(a), in a controlled convex setting, we observe that FedProx consistently interpolates between GlobalTrain and LocalTrain. As $\lambda$ increases, FedProx transitions from behaving similarly to LocalTrain towards GlobalTrain. When statistical heterogeneity is low, both GlobalTrain and FedProx with larger $\lambda$ values achieve better statistical performance. However, as heterogeneity increases, LocalTrain and FedProx with smaller $\lambda$ values show better results. The figure also demonstrates that with an adaptive choice of $\lambda$ outlined in Theorem 1, FedProx consistently achieves a low error across varying levels of heterogeneity. See more results and discussion in Appendix C. Note that as statistical heterogeneity increases, LocalTrain experiences a slight rise in error due to the increased difficulty of the statistical problem. However, it remains the least affected by heterogeneity since it relies solely on local data. In Figure 1(b) and 1(c), for the MNIST dataset, which includes both convex logistic regression and non-convex CNN settings, we observe similar patterns. FedProx again strikes a balance between GlobalTrain and LocalTrain, adapting its performance as $\lambda$ changes. While GlobalTrain performs best in this scenario due to the relatively weak statistical heterogeneity, FedProx, with increasing $\lambda$, moves closer to GlobalTrain performance. In Figure 4 in Appendix C, we also evaluate the local model’s statistical error and observe that as statistical heterogeneity increases, FedProx with smaller $\lambda$ values can surpass GlobalTrain.
Figure 2: Effect of local updates on FedProx performance. Smaller $\lambda$ reduces communication with more local updates, while larger $\lambda$ benefits marginally from increasing local updates and requires more frequent communication.
6.3 Optimization Aspect: Communication-Computation Trade-off
In this section, we explore the trade-off between communication and computation in FedProx, controlled by varying the personalization degree.
Communication Cost. In the bottom row of Figure 1, we show the change in the total error of the algorithm (i.e., the difference of the model in $t$th round between the true underlying model) over communication rounds. The first observation is that as $\lambda$ increases, the convergence rate w.r.t. communication rounds slows across all settings. FedProx with a large $\lambda$ (closer to GlobalTrain) needs the most communication rounds to converge, while smaller $\lambda$ values (closer to LocalTrain) result in faster convergence. This result aligns with our theory (c.f. Corollary 1), showing that Fedprox with less personalization relies more heavily on communication. On the statistical side, Figure 1(d) shows that when statistical heterogeneity is low, larger $\lambda$ leads to the smallest total error. As $R$ increases, smaller $\lambda$ (more personalization) results in better performance.
Computation Cost. In Figure 2, we investigate how the number of local updates per communication round impacts the performance of FedProx at different personalization degree. Additional results in the non-convex setting can be found in Appendix C. Notice that for models with smaller $\lambda$ (higher personalization), increasing the number of local updates significantly improves convergence, making the algorithm less dependent on frequent communication. This aligns with our analysis: more personalized models rely less on communication and benefit more from local computation.
7 Conclusion
In this paper, we analyzed the effect of personalization in FedProx. With an adaptive choice of personalization degree, FedProx achieves a nearly minimax-optimal statistical accuracy under different statistical heterogeneity degrees and consistently outperforms pure local training. As a minor contribution, we provably show that personalization reduces communication costs without increasing the computation overhead. The results are validated through extensive experiments.